# VideoAgent: Self-Improving Video Generation

### v2

@@ 4.2 End-to-End Task Success
VideoAgent-Online (Iter1) 41.3% 32.0% 17.3% 12.0% 45.3% 14.7%
VideoAgent-Online (Iter2) 44.0% 29.3% 18.7% 18.7% 46.7% 16.0%
VideoAgent-Online-Replan 80.0% 97.3% 40.0% 22.7% 72.0% 40.0%
@@
VideoAgent-Online (Iter1) 38.7% 13.3% 36.0% 0.0% 4.0% 23.2%
VideoAgent-Online (Iter2) 49.3% 21.3% 44.0% 1.3% 1.3% 26.4%
VideoAgent-Online-Replan 58.7% 36.0% 85.3% 8.0% 10.7% 50.0%
@@
We report the task success of baselines and VideoAgent in Table 1. Following Ko et al. (2023), we measure the average success across 3 camera poses with 25 seeds for each camera pose. Without online environment access, VideoAgent already improves the overall success rate from self-conditioned consistency alone over the baseline (from 19.6% to 22.3%). Furthermore, some specific tasks such as faucet-close shows drastic improvement from 12% to 46.7%. With online data collection, VideoAgent-Online further improves the success rate, and each online iteration (rolling out the policy, collecting successful trajectories, and continuing finetuning the policy) further improves the performance. The benefit of VideoAgent persists when replanning is introduced, brining the overall task success to 50%, achieving state-of-the-art in this setup. Qualitative improvement in the refined videos can be found in Figure 8 in Appendix H,
@@
4.3 Understanding the Effect of Different Components in VideoAgent
@@
4.3.1 Effect of Different VLM Feedback.
@@
In the previous section, we only used VLM during inference to determine when to stop refining a generated video. However, it is natural to wonder if information-rich feedback from the VLM, such as language descriptions of which part of a generated video to improve, might lead to better refined videos. To answer this question, we propose a few variants of VideoAgent according to the feedback available when training the video refinement model as in Equation 10. Specifically, we use VideoAgent to denote training the video refinement model only conditioned on the original task description. VideoAgent-Binary denotes additionally conditioning on whether a generated video is determined to be successful by the VLM. VideoAgent-Suggestive denotes conditioning additionally on language feedback from the VLM on which part of the video needs improvement and how the video can be improved. We train these three versions of the video refinement model, and report the overall task success from Meta-World in Table 3. We see that VideoAgent-Binary improves upon the base VideoAgent, while training with descriptive feedback in VideoAgent-Suggestive leads to even better performance. This suggests that richer feedback from the VLM can facilitate better training of the video refinement model. Improvement for each individual task can be found in the Appendix F.
4.3.2 Effect of Refinement and Online Iterations.
Next, we want to understand whether more refinement iterations and online finetuning iterations generally lead to higher task success. We found that while different tasks require a different number of iterations to achieve the best performance, VideoAgent does perform better as the number of refinement and online iterations increases, as shown in Figure 4 and Figure 4. During video refinement, specific tasks such as handle-press and faucet-close continue to see improvement even at the fifth refinement iteration. Faucet-close especially benefits from more refinement iterations, bringing success rate from 17.3% to 49.3% after five refinement iterations. The improved task success rates across refinement and online iterations suggests that self-conditioning consistency discussed in Section 3.1 and online interaction discussed in Section 3.3 can indeed effectively reduce hallucination and improve physical plausibility in the generated videos.
@@
4.3.3 Accuracy of VLM feedback on Generated Videos.
@@
Since this work is among the first to leverage a VLM to give feedback for video generation, it is crucial to understand whether a VLM can in fact achieve a reasonable accuracy in providing feedback for video generation. To quantify the performance of a VLM, we use human labels on whether a generated video is acceptable as the ground truth, and measure precision, recall, F1-score, and accuracy based on whether GPT-4 Turbo thinks the generated video is acceptable according to trajectory smoothness (consistent across sequential frames), physical stablility, and achieving the goal (See full prompt in Appendix A). We report the average result across 36 generated videos from the Meta-World dataset in Table 4. We see that the original prompt we used (Unweighted) achieves 69% accuracy, suggesting that the VLM is capable of judging generated videos. Since VideoAgent uses multiple refinement iterations, we want to avoid false positives where a bad video is accidentally accepted. We can achieve this by penalizing false positives through reweighting its cost in the prompt, which leads to the VLM rejecting videos when the VLM is uncertain about the video’s acceptability. This adjustment results in a significant increase in precision as shown in Table 4. This weighted version of the prompt is used in the experiments in Section 4.2.
@@
In the AVDC experimental setup, center cropping the third camera (what is used in the pipeline) often results in most of the robot arm being outside of the frame. We found that the accuracy of the VLM is affected by such partial observatbility. As shown in Table 4, removing the third camera from the prompt leads to much higher accuracy.
@@
While the VLM can provide binary feedback on whether a generated video is acceptable, we also measure the accuracy of the VLM in giving more descriptive feedback such as identifying the issue and providing suggestions on how to improve the video. We use three examples with human written language feedback as prompt for in-context learning. GPT-4 Turbo achieves 73.5% accuracy on identification and 86.1% accuracy on suggestion, as evaluated by humans. This result is highly encouraging and opens up future directions of leveraging descriptive feedback from VLMs to improve video generation.
@@ 4.4 Evaluating Self-Refinement on Real-World Videos
Video Score Visual Quality 1.97 ± 0.003 2.01 ± 0.003
@@
We report the average across 2250 videos generated from the AVDC baseline and from VideoAgent in Table 5. VideoAgent performs better according to all metrics except for Dynamic Degree from Video-Score (which shows similar performance between the two methods). Notably, the gain is significant in metrics critical for real-world videos, such as CLIP Score, Factual Consistency, and Text-to-Video Alignment. Improvement in Flow Consistency and Temporal Consistency suggests that VideoAgent produces smoother and more physically plausible videos that adhere better to the physical constraints of the real-world. This directly translates to better performance in real-world robotic tasks in Table 1.
@@
Next, we qualitatively evaluate generated videos from the AVDC baseline and from VideoAgent. We collect 50 generated videos from each model and conduct human evaluation on whether a generated video looks realistic. Videos with refinement from VideoAgent improves the acceptance rate by 22% as shown in Table 5. We further show an example video with and without refinement in Figure 5, where the baseline (middle row) hallucinates (the bowl disappears) whereas VideoAgent produces the video that completes the task (bottom row). We also present a more fine-grained analysis of Visual Quality, Temporal Consistency, Dynamic Degree, Text to Video Alignment, and Factual Consistency evaluated by humans in the Appendix G with the metrics in Table 8, which further echos the results of human evaluations presented in Table 5.
5 Related Work
Feedback and Self-improvement in LLMs.
Incorporating feedback and preference signals from feedback into the finetuning process of LLMs, has led to the enormous popularity and practical usability of the current versions of LLMs as chatbots  (Casper et al., 2023). Preference feedback from humans or other AI systems  (Ouyang et al., 2022a; Lee et al., 2023; Kaufmann et al., 2023) are first collected to train a reward model to guide the LLM’s generation or do implicit policy optimization  (Schulman et al., 2017; Rafailov et al., 2024). Furthermore LLMs have shown the ability to further improve by iterative refinement during finetuning and inference  (Zelikman et al., 2022; Yuan et al., 2024; Tian et al., 2024). We incorporate this reward driven improvement mechanism in our work, but unlike the LLM setting where the feedback came from a reward model or some proxy of this prefernce model, in our VideoAgent we use natural feedback from real world when simulated videos are turned into actions that are executed in the real world.
Image and Video Generation and Editing.
With the advent of large scale foundation models pretrained on internet scale data  (Bommasani et al., 2021), generation of super realistic multimodal content has become easier. Text generation, image or video generation, and cross-modal generation  (OpenAI et al., 2024; Reid et al., 2024; Wu et al., 2021; Ho et al., 2022; Singer et al., 2022; Yang et al., 2023a; Blattmann et al., 2023) has seen major advancements leveraging the autoregressive and diffusion based models architectures. And moving beyond simple generation, these models have been leveraged for guided text, image or video editing and enhancement  (Huang et al., 2024) to improve textual and visual aesthetics applied mostly to generative media  (Zhang et al., 2023). But none of these existing methods focus on grounding a generative simulator in the real world to perform more complex interactive multi-turn agentic and physical tasks needing both perception and control. To solve this bottleneck, we propose VideoAgent to self-improve or edit generated plan based on grounded feedback from real-world to execute robot manipulation tasks.
Video Generation for Robot Learning.
Video-based learning for robotics has been extensively studied  (Nair et al., 2022; Bahl et al., 2022; Shao et al., 2021; Chen et al., 2021; Pari et al., 2022; Sharma et al., 2019; Sun et al., 2018; Lee & Ryoo, 2017). Methods use video datasets for visual representation learning, goal extraction, and dynamic models for planning  (Finn & Levine, 2017; Kurutach et al., 2018), or imitation learning from expert actions  (Fang et al., 2019; Wang et al., 2023; Mani et al., 2024), state sequences  (Torabi et al., 2019; Lee et al., 2021; Karnan et al., 2022), and pretraining on videos followed by RL  (Baker et al., 2022; Escontrela et al., 2023). Recently, generative models have advanced video-based learning and planning, framing decision-making as text-conditioned video generation to predict trajectories  (Du et al., 2024; Ko et al., 2023; Wen et al., 2023). Vision-language and text-to-video models generate long-horizon plans for robotic tasks through abstract, visually grounded planning  (Du et al., 2023; Ajay et al., 2024). Generative models also simulate agent-environment interactions, enabling zero-shot deployment  (Yang et al., 2023b), and test-time feedback for replanning  (Bu et al., 2024). Unlike these, our VideoAgent improves video generation during training with real-world feedback and refines actions through test-time self-iteration and replanning.
@@ 6 Conclusion and Future Work
7 Acknowledgement
Thanks to Doina Precup and Douglas Eck for reviewing draft versions of this manuscript. Thanks to Wenhu Chen for his inputs on VLM feedback. We gratefully acknowledge Sebastian Fischmeister for providing funds for this work, and the Chippie cluster at the University of Waterloo for additional GPU resources.
Appendix A Prompt Structure for VLM Feedback
A.1 Binary Classification
@@ A.2 Identification and Suggestion:
Appendix B Task Descriptions and In-Context Examples for VLM Feedback
@@
Appendix C Dataset Descriptions in Detail
@@
Appendix D Extended Experiments
D.1 Further Analysis of VideoAgent-Online
We train VideoAgent-Online for multiple iterations and observe that after 2 iterations, the results start to stabilize. The results for iteration 3 are shown in table 6.
@@ D.1 Further Analysis of VideoAgent-Online
VideoAgent-Online(Iter1) 41.3% 32.0% 17.3% 12.0% 45.3% 14.7%
VideoAgent-Online(Iter2) 44.0% 29.3% 18.7% 18.7% 46.7% 16.0%
VideoAgent-Online(Iter3) 46.7% 28.0% 18.7% 18.7% 45.3% 20.0%
@@
VideoAgent 46.7% 12.0% 36.0% 0.00% 1.3% 22.3%
VideoAgent-Online(Iter1) 38.7% 13.3% 36.0% 0.00% 4.0% 23.15%
VideoAgent-Online(Iter2) 49.3% 21.3% 44.0% 1.33% 1.33% 26.4%
VideoAgent-Online(Iter3) 48.0% 21.3% 42.0% 1.33% 1.33% 26.5%
Table 6: Meta-World Result. The mean success rates of VideoAgent combined with Online and Replan modules as compared to the AVDC baseline
Appendix E Architectural Details of VideoAgent
E.1 Video Diffusion training details
Appendix F VLM Feedback for Correction

### v3

@@ 3.1 Refinement through self-conditioning consistency
VideoAgent-Online (Iter1) 48.0% 40.0% 24.0% 12.0% 42.7% 36.0%
VideoAgent-Online (Iter2) 58.7% 50.7% 28.0% 18.7% 53.3% 41.3%
VideoAgent-Online-Replan 82.7% 97.3% 40.0% 26.7% 73.3% 44.0%
@@
VideoAgent-Online (Iter1) 53.3% 28.0% 52.0% 1.3% 5.3% 31.2%
VideoAgent-Online (Iter2) 58.7% 36.0% 64.0% 1.3% 9.3% 38.2%
VideoAgent-Online-Replan 74.7% 46.7% 86.7% 8.0% 10.7% 53.7%
3.3 Self-improvement through online finetuning
In addition to video refinement through self-conditioning consistency and inference-time compute as described in Section 3.1 and Section 3.2, we can further characterize the combination of video generation and video refinement as a policy, which can be improved by training on additional data collected from the environment during online interaction. Specifically, the goal is to maximize the expected returns of a policy through trial-and-error interaction with the environment:
where $\mathcal{R}$ is the true reward function, $\mathcal{E}$ is the interactive environment, and $\pi_{\theta}$ corresponds to the combination of $f_{\theta}$ and $\hat{f}_{\theta}$.
where $\rho$ is the optical flow model that maps the refined video to low-level control actions. See Algorithm 3 in Appendix A for details of online policy finetuning.
4 Experiments
We now evaluate the performance of VideoAgent. We introduce the experimental settings and variants of VideoAgent in Section 4.1, measure end-to-end success rate of VideoAgent against the baselines in Section 4.2, and study the effect of different components of VideoAgent in Section 4.3. Finally, we show that VideoAgent is effective in improving the quality of real robotic videos in Section 4.4.
4.1 Datasets and Experimental Setups
Datasets and Environments.
We follow the same evaluation setting as Ko et al. (2023), which considers three datasets: Meta-World (Yu et al., 2020), iTHOR (Kolve et al., 2017), and BridgeData V2 (Walke et al., 2023). Meta-World consists of 11 robotic manipulation tasks performed by a simulated Sawyer arm, with video demonstrations captured from three distinct camera angles. iTHOR is a simulated 2D object navigation benchmark, where an agent searches for specified objects across four room types. BridgeData V2 is a real-world dataset of robotic manipulation. See more details of datasets and environments in Appendix D.
Baselines and VideoAgent Variants.
We consider the following methods for comparison:
  • AVDC (baseline). This is the Actions from Video Dense Correspondences (Ko et al., 2023) baseline, which synthesizes a video and predicts optical flow to infer actions.
  • VideoAgent-Replan. This variant incorporates online filtering of successful trajectories with the replanning mechanism, where replanning is conducted first, and more successful trajectories after replanning are added back to the training data.
4.2 End-to-End Task Success
@@
We report the task success rates of baselines and VideoAgent in Table 1. Following (Ko et al., 2023), we evaluate performance across three camera poses with 25 seeds per pose. Without online environment access, VideoAgent improves the overall success rate through self-conditioning consistency alone from 19.6% (AVDC) to 22.3%. For certain difficult tasks, e.g., faucet-close, VideoAgent improves performance from 12.0% to 46.7%. With online data collection, VideoAgent-Online further improves success rates with each additional online iteration of rolling out the policy, collecting successful trajectories, and finetuning. VideoAgent-Online can be further combined with replanning, achieving 53.7% overall success, surpassing prior state-of-the-art on this benchmark. Detailed baseline comparisons are provided in Appendix E.2, and qualitative improvements in refined videos are shown in Figure 9 in Appendix I.
@@
4.3 Effect of Different Components in VideoAgent
@@
4.3.1 Effect of Different VLM Feedback
@@
In the previous section, we only used VLM during inference to determine when to stop refining a generated video. However, it is natural to wonder if information-rich feedback from the VLM, such as language descriptions of which part of a generated video to improve, might lead to better refined videos. To answer this question, we propose a few variants of VideoAgent according to the feedback available when training the video refinement model as in Equation 10. Specifically, we use VideoAgent to denote training the video refinement model only conditioned on the original task description. VideoAgent-Binary denotes additionally conditioning on whether a generated video is determined to be successful by the VLM. VideoAgent-Suggestive denotes conditioning additionally on language feedback from the VLM on which part of the video needs improvement and how the video can be improved. We train these three versions of the video refinement model, and report the overall task success from Meta-World in Table 3. We see that VideoAgent-Binary improves upon the base VideoAgent, while training with descriptive feedback in VideoAgent-Suggestive leads to even better performance. This suggests that richer feedback from the VLM can facilitate better training of the video refinement model. Improvement for each individual task can be found in the Appendix G.
4.3.2 Effect of the Number of Iterations.
Next, we want to understand whether more refinement iterations and online finetuning iterations lead to higher task success. We found that while different tasks require a different number of iterations to achieve the best performance, VideoAgent does perform better as the number of refinement and online iterations increases, as shown in Figure 4 and Figure 4. During video refinement, specific tasks such as handle-press and faucet-close continue to show improvement even at the fifth refinement iteration. Faucet-close especially benefits from more refinement iterations, bringing success rate from 24.0% to 58.7% after five refinement iterations. The improved task success rates across refinement and online iterations suggests that self-conditioning consistency discussed in Section 3.1 and online interaction discussed in Section 3.3 can indeed effectively reduce hallucination and improve physical plausibility in the generated videos.
@@
4.3.3 Accuracy of VLM feedback
@@
Since this work is among the first to leverage a VLM to give feedback for video generation, it is crucial to understand whether a VLM can in fact achieve a reasonable accuracy in providing feedback for video generation. To quantify the performance of a VLM, we use human labels on whether a generated video is acceptable as the ground truth, and measure precision, recall, F1-score, and accuracy based on whether GPT-4 Turbo thinks the generated video is acceptable according to trajectory smoothness (consistent across sequential frames), physical stability, and achieving the goal (See full prompt in Appendix B). We report the average result across 36 generated videos from the Meta-World dataset in Table 4. We see that the original prompt we used (Unweighted) achieves 69% accuracy, suggesting that the VLM is able to somewhat judge generated videos but not always accurately. Since VideoAgent uses multiple refinement iterations, we want to avoid false positives where a bad video is accidentally accepted. We can achieve this by penalizing false positives through reweighing its cost in the prompt, which leads to the VLM rejecting videos when the VLM is uncertain about the video’s acceptability. This adjustment results in a significant increase in precision as shown in Table 4. This weighted version of the prompt is used in the experiments in Section 4.2.
@@
In the AVDC experimental setup, center cropping the third camera (what is used in the pipeline) often results in most of the robot arm being outside of the frame. We found that the accuracy of the VLM is affected by such partial observability. As shown in Table 4, removing the third camera from the prompt leads to much higher accuracy.
@@
While VLM can provide binary feedback on whether a generated video is acceptable, we also measure the accuracy of the VLM in giving more descriptive feedback such as identifying the issue and providing suggestions on how to improve the video. We use three examples with human written language feedback as prompt for in-context learning. GPT-4 Turbo achieves 73.5% accuracy on identification and 86.1% accuracy on suggestion, as evaluated by humans. This result is highly encouraging and opens up future directions of leveraging descriptive feedback from VLMs to improve video generation.
@@ 4.4 Evaluating Self-Refinement on Real-World Videos
Visual Quality 1.97 ± 0.003 2.01 ± 0.003
@@
We report the average across 2250 videos generated from the AVDC baseline and from VideoAgent in Table 5. VideoAgent performs better according to all metrics except for Dynamic Degree from Video-Score (which shows similar performance between the two methods). Notably, the gain is significant in metrics critical for instruction following and real-world videos, such as CLIP Score, Factual Consistency, and Text-to-Video Alignment. Improvement in Flow Consistency and Temporal Consistency suggests that VideoAgent produces smoother and more physically plausible videos that adhere better to the physical constraints of the real-world. This directly translates to better performance in real-world robotic tasks in Table 1.
@@
Next, we qualitatively evaluate generated videos from the AVDC baseline and from VideoAgent. We collect 50 generated videos from each model and conduct human evaluation on whether a generated video looks realistic. Videos with refinement from VideoAgent improves the acceptance rate by 22% as shown in Table 5. We further show an example video with and without refinement in Figure 5, where the baseline (middle row) hallucinates (the bowl disappears) whereas VideoAgent produces the video that completes the task (bottom row). We also present a more fine-grained analysis of Visual Quality, Temporal Consistency, Dynamic Degree, Text to Video Alignment, and Factual Consistency evaluated by humans in the Appendix H with the metrics in Table 9, which further echos the results of human evaluations presented in Table 5.
@@ 5 Conclusion and Future Work
  • VideoAgent trades off inference-time compute for better performance by iteratively refines generated video plans under the guidance of a VLM. Exploring other inference-time search strategies to further improve video sample quality is an interesting area of research.
6 Impact Statement
VideoAgent introduces a novel self-conditioning consistency mechanism that enables iterative refinement of generated video plans, significantly improving long-horizon task completion. By leveraging previously generated video segments for refinement, VideoAgent mitigates hallucinations and enhances temporal consistency without requiring extensive interaction with the environment. This reduces the need for costly and time-consuming data collection while still achieving state-of-the-art success rates in simulated robotic environments. Furthermore, VideoAgent ’s ability to refine plans without relying on replanning makes it highly adaptable to real-world applications, including robotics, autonomous systems, and video-based reinforcement learning. This work advances scalable and generalizable video generation techniques, contributing to the broader goal of AI agents that can reason and act through visual understanding.
Appendix
Appendix A Algorithms
for iteration $=1$ to $N$ do
       Sample $\{(\mathbf{x}^{(0)},g)\}\sim\mathcal{D}$ and $t\sim\text{Uniform}(\{0,1,\dots,T\})$  Compute vanilla diffusion loss:
$\mathcal{L}_{\text{video-diffusion}}=\left\|f_{\theta}(\mathbf{x}^{(t)},t)-\mathbf{x}^{(0)}\right\|^{2}$  Generate $\hat{\mathbf{x}}$ following Equation 11 and sample $\texttt{feedback}\sim\hat{\mathcal{R}}(\cdot|\hat{\mathbf{x}})$  Compute consistency loss:
$\mathcal{L}_{\text{SC-consistency}}=\left\|\hat{f}_{\theta}(\hat{\mathbf{x}},\mathbf{x}^{(t)},t\,|\,\texttt{feedback})-\mathbf{x}^{(0)}\right\|^{2}$  Update parameters:
$\theta\leftarrow\theta-\gamma\nabla_{\theta}\left(\mathcal{L}_{\text{video-diffusion}}+\mathcal{L}_{\text{SC-consistency}}\right)$
Algorithm 2 VLM Guided Replan
[…]
       $\hat{\mathbf{x}}\leftarrow\pi_{\theta}(x_{0},g)$for $i=0$ to max_refine_iterations do
[…]
for iteration $i=1$ to $N$ do
       $\mathcal{D}_{\text{new}}\leftarrow\emptyset$for each $(\cdot,g)$ in $\mathcal{D}$ do
             $x_{0}\leftarrow\mathcal{E}.\text{reset}(g)$$\hat{x}_{\text{refined}}\sim\pi_{\theta}(x_{0},g)$if $\mathcal{R}(\rho(\hat{x}_{\text{refined}}))$ then
                   $\mathcal{D}_{\text{new}}\leftarrow\mathcal{D}_{\text{new}}\cup(\hat{x}_{\text{refined}},g)$
      $\mathcal{D}\leftarrow\mathcal{D}\cup\mathcal{D}_{\text{new}}$  Finetune $\theta$ using Algorithm 1 on $\mathcal{D}$
Appendix B Prompt Structure for VLM Feedback
B.1 Binary Classification
@@ B.2 Identification and Suggestion:
Appendix C Task Descriptions and In-Context Examples for VLM Feedback
@@
Appendix D Dataset Descriptions in Detail
@@
D.1 Additional trajectories per iteration during online training
We collect 15 successful trajectories for each task during every iteration. This standardization helps address task imbalance, as task success rates are higher for certain tasks compared to others. By ensuring a fixed number of successful trajectories per task, we prevent overfitting to easier tasks and maintain balanced model performance across the entire task set.
Appendix E Extended Experiments
E.1 Videos to action conversion
We employ the GMFlow optical flow model to predict dense pixel movements across frames. These predicted flows serve as the foundation for reconstructing both object movements and robot motions depicted in the video. The flow predictions allow us to interpret the temporal evolution of the video in terms of actionable physical dynamics. The optical flow essentially provides a dense correspondence of pixel movements between consecutive frames, which is then used to infer the relative motion of objects and the robot. This mapping bridges the gap between the high-dimensional video representation and the low-level control commands required to execute the tasks in a simulated or real environment.
This method ensures that the generated video plans are actionable and aligned with the task-specific dynamics, making the video generation process directly relevant to downstream policy learning and execution.
E.2 Baseline experiments on Metaworld
We conduct experiments on additional baselines including, Behavioral Cloning (BC), UniPi (with replan), VLP and Diffusion policy. Table 6 consists of these results.
@@ E.2 Baseline experiments on Metaworld
BC-Scratch 21.3% 36.0% 0.0% 0.0% 34.7% 12.0%
BC-R3M 1.3% 58.7% 0.0% 0.0% 36.0% 4.0%
UniPi (with Replan) 0.0% 36.0% 0.0% 0.0% 6.7% 0.0%
AVDC 30.7% 28.0% 21.3% 8.0% 34.7% 17.3%
VLP 33.3% 28.0% 17.3% 8.0% 36.0% 18.7%
Diffusion Policy 45.3% 45.3% 8.0% 0.0% 40.0% 18.7%
AVDC-Replan 72.0% 89.3% 37.3% 18.7% 60.0% 24.0%
@@
VideoAgent (Iter2) 48.0% 40.0% 24.0% 12.0% 42.7% 36.0%
VideoAgent (Iter3) 58.7% 50.7% 28.0% 18.7% 53.3% 41.3%
VideoAgent-Replan 82.7% 97.3% 40.0% 26.7% 73.3% 44.0%
@@
BC-Scratch 18.7% 22.7% 28.0% 0.0% 0.0% 15.4%
BC-R3M 18.7% 17.3% 37.3% 0.0% 1.3% 16.2%
UniPi (with Replan) 4.0% 9.3% 13.3% 4.0% 0.0% 6.1%
AVDC 12.0% 17.3% 41.3% 0.0% 5.3% 19.6%
VLP 30.7% 10.7% 33.3% 0.0% 1.3% 19.8%
Diffusion Policy 22.7% 58.7% 21.3% 4.0% 1.3% 24.1%
AVDC-Replan 53.3% 24.0% 81.3% 8.0% 6.7% 43.1%
VideoAgent 46.7% 12.0% 36.0% 0.0% 1.3% 22.3%
VideoAgent (Iter2) 53.3% 28.0% 52.0% 1.3% 5.3% 31.2%
VideoAgent (Iter3) 58.7% 36.0% 64.0% 1.3% 9.3% 38.2%
VideoAgent-Replan 74.7% 46.7% 86.7% 8.0% 10.7% 53.7%
E.3 Further Analysis of VideoAgent-Online
We train VideoAgent-Online for multiple iterations and observe that after 2 iterations, the results start to stabilize. The extra results for iteration 3 are also shown in table 7.